Grade 11General Physics

Q.Some persons hire a boat for 4 hours.The river flows at a speed of 5km/hr and boat moves at a speed of 15km/hr relative to the water.How far along the flow direction,they can go if they have to return in 4hrs?
Ans)80/3 km

jiten
8 Years agoGrade 11

1 Answer

Khimraj
8 Years ago
Let that distance is x km
x/(15 +5) + x/(15-5) = 4
x/20 + x/10 = 4
3x/20 = 4
then x = 80/3 km.
Hope it clears. If you like answer then please approve it.